### v1

# REvolve: Reward Evolution with Large Language Models for Autonomous Driving

Abstract
Designing effective reward functions is crucial to training reinforcement learning (RL) algorithms. However, this design is non-trivial, even for domain experts, due to the subjective nature of certain tasks that are hard to quantify explicitly. In recent works, large language models (LLMs) have been used for reward generation from natural language task descriptions, leveraging their extensive instruction tuning and commonsense understanding of human behavior. In this work, we hypothesize that LLMs, guided by human feedback, can be used to formulate human-aligned reward functions. Specifically, we study this in the challenging setting of autonomous driving (AD), wherein notions of "good" driving are tacit and hard to quantify. To this end, we introduce REvolve, an evolutionary framework that uses LLMs for reward design in AD. REvolve creates and refines reward functions by utilizing human feedback to guide the evolution process, effectively translating implicit human knowledge into explicit reward functions for training (deep) RL agents. We demonstrate that agents trained on REvolve-designed rewards align closely with human driving standards, thereby outperforming other state-of-the-art baselines.
11footnotetext: Equal Contribution.
1 Introduction
[…]
Recently, large language models (LLMs) have shown impressive abilities in instruction following [29, 38], search and black-box optimization[28, 62, 65], self-reflection [32, 49], and code generation [6, 42]. Owing to their vast training and world knowledge [17, 19, 59], they can readily incorporate commonsense priors of human behavior. Together, these abilities make LLMs an ideal candidate for reward generation from natural language (NL) task descriptions and refining them using language-based feedback loops. Indeed, recent works like Language to Rewards [63] and Text2Reward [61] have utilized LLMs for this purpose.
In particular, Eureka [30] utilizes LLMs to design dense and interpretable reward functions, achieving improved performance in dexterous manipulation tasks. It does so by generating multiple reward functions and greedily iterating with the best candidate over multiple iterations. We identify this greediness as a major limitation as it risks premature convergence and leads to a waste of resources (by preemptively discarding weaker reward functions). Moreover, Eureka relies on having access to fitness functions in order to assess the quality of learned behaviors. This is particularly challenging for complex tasks, such as AD, which require human evaluations.
To this end, we introduce Reward Evolve (REvolve), a novel evolutionary framework using LLMs, specifically GPT-4, to output reward functions (as executable Python codes) for AD and evolve them based on human feedback. In particular, REvolve offers several key advantages:
Figure 1: An overview of REvolve. Given the task of AD and abstracted environment variables, a reward designer $G$ (LLM) outputs a population of reward functions $R\in\mathcal{R}$, each used to train an AD policy $\pi(R)$ in driving simulation. Then, we collect human preferences and natural language (NL) feedback on pairs of policy rollouts $\theta\sim\Theta_{\pi(R)}$ through a human user feedback interface. Policy (and thus, corresponding reward function) fitness $\sigma$ is calculated, and the fittest individuals, along with their NL feedback $\lambda$, are refined by $G$. The process leverages genetic programming for evolution. The flames symbolize trainable parameters.
(1) Framing reward design as a search problem. Based on evolutionary algorithms [2], REvolve evolves a population of reward functions by applying genetic operations such as mutation, crossover, and migration. This enhances genetic diversity and prevents premature convergence in complex search space (in comparison to greedy search), all without incurring additional computational costs compared to Eureka.
(2) Utilizing human feedback to guide the search. Human preference data is directly mapped into fitness scores, effectively allowing humans to serve as fitness functions. Additionally, the framework incorporates qualitative natural language feedback, offering GPT-4 high-level insights into which aspects to prioritize while refining the reward functions. Together, this ensures that the reward functions – and thus the learned policies – align with human driving standards. Unlike inverse reinforcement learning (IRL) approaches, which require extensive demonstrations [46] or interventions [45] to infer human behavior indirectly, REvolve uses minimal feedback to directly refine reward functions.
(3) Eliminating the need for additional reward model training. Unlike reinforcement learning with human feedback (RLHF) [9], which relies on a black-box reward model (trained on human preferences between outputs) to align LLMs with goals like safety and reliability [36], REvolve requires no extra models. REvolve-designed reward functions are interpretable (as executable Python code) compared to RLHF rewards, making it ideal for safety-critical applications, such as AD [8].
Figure 1, provides an overview of REvolve for AD within a simulated environment.
2 Problem Formalization
Let us first formally introduce the reward design problem (RDP) [51], given by the tuple $\langle M,\mathcal{R},\pi,F\rangle$, where $M=(S,A,T)$ is the world model consisting of state space $S$, action space $A$, and transition function $T$. $\mathcal{R}$ is the set of all reward functions where each $R\in\mathcal{R}$ is a reward function that maps a state-action pair to a real-valued scalar $R:S\times A\rightarrow\mathbb{R}$. The tuple $(M,R)$ constitutes a Markov decision process [53]. Furthermore, $\pi:S\rightarrow\Delta A$ is a policy that maps a state to a distribution over actions, which is learned using a reward function $R$.
Moreover, the RDP includes a function $F$ that evaluates the fitness (i.e. quality) of the trained policies. Traditionally, $F$ maps policy rollouts, i.e. rollouts of agent-environment interactions, to real-valued scalars. Formally, we denote this by $F:\Omega(\Theta_{\pi})\rightarrow\mathbb{R}$, where $\Theta_{\pi}$ is the random variable modeling the probability distribution of rollouts induced by $\pi$ and $\Omega(\Theta_{\pi})$ is its sample space.
For REvolve, we generalize this to also include natural language feedback. Specifically, we redefine $F:\Omega(\Theta_{\pi})\rightarrow\mathbb{R}\times\mathcal{L}$, where $\mathcal{L}$ is the set of all strings (including natural language feedback). Note that the fitness function is derived from human evaluations of the trained policies, thereby rendering it non-differentiable.
In order to instantiate the RDP tuple for REvolve, we assume that we are given a reward designer $G:\mathcal{L}\rightarrow\mathcal{R}$ that maps from the set of all strings to the set of all reward functions $\mathcal{R}$. The objective now is to find a reward function $R\in\mathcal{R}$ that maximizes the expected fitness score of the policy $\pi$ trained on $R$, which we denote by $\pi(R)$. Formally we express the optimization problem as:
$R^{*}\coloneqq\operatorname*{arg\,max}_{R\in\mathcal{R}}\Bigl{[}\mathbb{E}_{\theta\sim\Theta_{\pi(R)}}\left[F(\theta)\right]\Bigr{]}$ (1)
Here, $\theta$ is a trajectory sampled from $\Theta_{\pi(R)}$. We also define the optimal policy $\pi^{*}$ as the policy trained with reward function $R^{*}$ given as $\pi^{*}\coloneqq\pi(R^{*})$.
An implicit assumption for reward design is the knowledge of observation and action variables in the environment. For real-world problems like AD, the assumption is supported by access to sensor data (such as speed from speedometers and distance covered from odometers), or by employing advanced vision-based models [7, 22, 64].
3 Reward Evolution
As the fitness function $F$ is based on implicit human understanding, it cannot be explicitly formulated and is therefore not amenable to gradient-based optimization. Therefore, we solve the RDP using a meta-heuristic approach through genetic programming (GP) [26]. Concretely, REvolve uses GP to output and iteratively refine a population of reward functions (individuals) by applying genetic operations such as mutation, crossover, and migration. In each iteration, we obtain new individuals, which are added to the population based on their fitness. This fitness is assessed through human evaluations of the policies trained with these reward functions – thus, guiding the evolutionary search.
Generally, GPs require significant domain expertise and rely on random initialization and perturbations for exploration, leading to scalability issues [37]. In contrast, LLMs can suggest changes that are intelligent and better aligned with the feedback [13]. Therefore, we combine the evolutionary search of REvolve with LLMs by using them as reward designers $G$ in the RDP.
We follow the island model of evolution [4] where we maintain multiple ($I$) sub-populations (or islands) in a reward database $D$. Island populations evolve by locally reproducing (mutation and crossover). Periodically, individuals migrate from one island to another. The island model helps enhance genetic diversity across the population and prevent premature convergence in complex search spaces [4]. Following existing works in reward design with LLMs [30, 61], we represent the reward functions as Python programs that seamlessly integrate with the GP framework. In what follows, we describe the four stages (initialization, reproduction, selection, termination) of our evolutionary framework. We give the formal description in Algorithm 1 – for brevity, we use $\pi$ instead of $\pi(R)$.
Algorithm 1 REvolve
1:Require fitness function $F$, reward designer $G$
2:Hyperparameters evolution generations $N$, number of sub-populations $I$, individuals per generation $K$, mutation probability $p_{m}$, crossover probability $p_{c}$
3:Input: Task description
4:# Initialization $\triangleright$ § 3.1
5:$D\coloneqq\{\big{(}R,\pi,\sigma,\lambda\big{)}_{i}\}_{i=1}^{K}$triangleright Initialize database
6:while not termination do
7:    # Reproduction $\triangleright$ § 3.2
8:    $D_{\text{temp}}\coloneqq\{\}$triangleright Initialize empty temporary database
9:    for $K$ individuals do
10:         is_mutate $\sim\texttt{Bernoulli}(p_{m})$triangleright Sample operation
11:         $P\sim\texttt{WeightedSample}(D,\{\sigma^{P}_{j}\}_{j=1}^{I})$triangleright Sample sub-population using average fitness
12:         if is_mutate then
13:             $R_{\text{new}}\coloneqq\texttt{Mutate}(G,D[P])$triangleright Mutate individual in sub-population
14:         else
15:             $R_{\text{new}}\coloneqq\texttt{Crossover}(G,D[P])$triangleright Crossover with $1-p_{m}$ probability
16:         end if
17:         $\pi\coloneqq\texttt{Train}(R_{\text{new}})$triangleright Train policy on new reward
18:         $D_{\text{temp}}\leftarrow\texttt{Update}(R_{\text{new}},P,\pi)$triangleright Add to temporary database
19:    end for
20:    # Selection $\triangleright$ § 3.3
21:    for $(R,P,\pi)\in D_{\text{temp}}$ do
22:         $\theta\sim\Theta_{\pi}$triangleright Sample rollouts from trained policy
23:         $(\sigma,\lambda)\coloneqq F(\theta)$triangleright Compute fitness from human feedback
24:         if $\sigma\geq\sigma^{P}$ then $\triangleright$ If individual fitness $\geq$ sub-population fitness
25:             $D[P]\leftarrow\texttt{Update}(R,\pi,\sigma,\lambda)$triangleright Add reward to database subpopulation
26:         end if
27:    end for
28:end while
29:# Termination $\triangleright$ § 3.4
30:$(R^{\ast},\pi^{\ast},\sigma^{\ast},\lambda^{\ast})\coloneqq\operatorname*{arg\,max}_{\sigma}D$triangleright Select tuple with the highest fitness score $\sigma$
31:Return $\pi^{\ast}$
3.1 Initialization
We start by creating a population of $K$ reward function individuals using GPT-4. This leverages GPT-4’s world knowledge [17] and its zero-shot instruction following abilities [38]. For each individual, we train a policy $\pi$, compute a fitness score $\sigma$, and obtain natural language feedback $\lambda\in\mathcal{L}$. The fitness score calculation and feedback collection are detailed in Section 3.3. Each individual and their respective policy, fitness score, and natural language feedback $(R,\pi,\sigma,\lambda)$ is then randomly assigned to a sub-population and stored in the reward database $D$.
3.2 Reproduction
Each successive generation of $K$ individuals is created by applying genetic operators (crossover and mutation), on the existing reward database individuals. We use GPT-4 as mutation and crossover operators with appropriate prompts. Due to space constraints, we provide prompt examples for mutation and crossover in Appendix D.
We begin by randomly selecting an operation – mutation with probability $p_{m}$ or crossover with probability $p_{c}=1-p_{m}$. Next, we sample a sub-population $P$ via weighted selection based on the average fitness scores $\sigma^{P}$ of each of the $I$ sub-populations. Subsequently, we sample parents (one for mutation, two for crossover) from $D[P]$ using weighted sampling based on individual fitness scores. These sampled individuals along with their natural language feedback $\lambda$ are used to prompt GPT-4 to create new individuals. These prompts direct GPT-4 to either modify a single reward component of the parent ($\texttt{Mutate}(G,D[P])$) or combine the best-performing reward components from two parents ($\texttt{Crossover}(G,D[P])$). Periodically, we enable the migration of individuals between sub-populations to enhance genetic diversity. Mutation and crossover operations are illustrated in Figure 2.
Figure 2: Illustration of how GPT-4 applies mutation and crossover to reward functions. Mutation (left): shows the modification of the "smoothness reward" component. A red ‘-’ sign indicates the line removed from the parent reward function, while a green ‘+’ sign indicates the line added to the new reward function. Crossover (right): demonstrates how parent reward functions are combined to create a child reward function, incorporating the most effective reward components from each parent.
3.3 Selection
Newly reproduced individuals (from the previous stage) are retained based on fitness scores, following a survival of the fittest approach. A fitness function $F$, tailored to the specific problem, evaluates the quality of each individual. The process involves (1) training a policy $\pi(R)$ until convergence, (2) sampling rollouts $\theta\sim\Theta_{\pi(R)}$, and (3) using the fitness function to obtain a tuple of (real-valued) fitness score and (string) NL feedback $(\sigma,\lambda)\coloneqq F(\theta)$. Individuals are retained if they contribute to an increase in the average fitness of their respective sub-populations $\sigma\geq\sigma^{P}$ where $P$ denotes the sub-population that the individual is added to [^1] To ensure genetic diversity, we add individuals based on surpassing the average and not maximum scores.. This ensures that the average fitness of the sub-population is monotonically increasing. In what follows, we detail the process of obtaining the tuple $(\sigma,\lambda)$ from the fitness function.
To compute the fitness score $\sigma$, we ask human evaluators to judge policy rollouts $\theta$ from different policies in a pairwise fashion. As shown by the "Human Feedback Interface" in Figure 1, these rollouts are presented as video segments $\approx$ 30-40 seconds long. The evaluators choose their preferred segment (ties are also possible). We then apply an Elo rating system [10], which transforms pairwise assessments into fitness scores $\sigma$. This avoids the need to learn a separate reward model, cf. RLHF [9]. Asking human evaluators to provide preferences instead of quantitative scores for each video has the advantage of mitigating the introduction of individual bias [9]. Details of the human feedback collection and Elo scoring are documented in Appendix B.1.
The NL feedback $\lambda$ about each pair of video segments is, likewise, collected through the same feedback interface. Evaluators comment (select checkboxes) on aspects of the driving they find satisfactory or need improvements. This feedback is subsequently integrated into the prompt for GPT-4 when reward function individuals are sampled for reproduction.
3.4 Termination
The evolutionary process repeats until it reaches a predetermined number of generations $N$, or when an individual attains a fitness score equivalent to human-level driving performance, as detailed in § 5.1. The final output is the fittest policy $\pi^{\ast}$ corresponding to the highest $\sigma$ in the database $D$. Note that $\pi^{\ast}$ is the best policy in the reward database according to human-assigned fitness scores $\sigma$, but it is not necessarily the optimal policy.
A minimal input-output example of the working of REvolve is shown in Appendix E.
4 Related Work
Reinforcement Learning and Human Feedback. Due to its complex and hard-to-specify goals [24], autonomous driving presents a challenging problem for reinforcement learning. Autonomous vehicles must maneuver through complex environments and interact with other road users, all the while strictly adhering to traffic regulations. Common techniques to overcome these challenges include learning from demonstrations (LfD) [46] or learning from intervention (LfI) [45]. However, they do not always ensure that the learned behaviors align with human values such as safety, robustness, and efficiency [15]. An alternative is integrating human feedback by using humans as the reward function in the learning process directly [23, 54]. However, this is rather resource-intensive as each rollout needs to be evaluated by a human. REvolve directly integrates human feedback (on the trained policy) into the reward design process. Instead of using humans as reward functions directly, human values, preferences, and knowledge can also be incorporated via so-called human-in-the-loop (HITL) RL. Several techniques exist, for instance, human assessment [9, 23, 54], human demonstration [46], and human intervention [45, 58]. They are particularly beneficial in scenarios like AD, where the reward function is challenging, and the nature of the desired outcome is complex. More recently, reinforcement learning from human feedback (RLHF) [9, 40] has been used successfully in training models like ChatGPT [36], providing guidance and alignment from user preferences. A further comparison of AD approaches with and human feedback is presented in Figure 3, which illustrates how REvolve uses human feedback to align learned policies with human driving standards without requirements for extra reward models, such as in the case of RLHF.
Figure 3: Comparison of REvolve with existing frameworks used to train AD agents. (a.) Expert Designed reward function $R$ to train the AD agent; (b.) learning from demonstrations (LfD) and learning from interventions (LfI), where the agent is trained to imitate the human; (c.) humans acting as reward functions within the training loop, by assessing policy rollouts and providing scalar rewards, which requires significant manual effort; (d.) Reinforcement learning from human feedback (RLHF), where human preference data is used to train an additional (black-box) reward model. This reward model is used as a proxy for human rewards to train the AD policy; (e) Proposed REvolve, which uses GPT-4 as a reward function generator $G$ and evolves them based on (minimal) human feedback on the rollouts sampled from the trained models. This feedback is directly incorporated into the reward design process. REvolve outputs interpretable reward functions, thereby avoiding learning an additional reward model. Here, $\pi\in\Pi$ is a trainable policy in the set of policies $\Pi$. The trainable blocks are denoted by the symbol of the flame.
Reward Design with LLMs. Recent works have explored LLMs for complex search and black-box optimization problems like hyperparameter optimization [28, 62, 65], reward design for RL [25, 30, 61], or robotic control [63]. Kwon et al. [25] use LLMs to directly output binary rewards by inferring user intentions in negotiation games, but this requires numerous queries to LLMs and limits the rewards to sparse binary values. Conversely, Language to Rewards (L2R) [63] and Text2Reward (T2R) [61] utilize LLMs to create dense and interpretable reward functions, improving them with human feedback. L2R, however, restricts its adaptability by employing manually crafted reward templates tailored to specific robot morphologies. In contrast, T2R outputs versatile reward functions in Python code. Eureka [30] further extends T2R by generating multiple reward functions and greedily iterating with the best reward function, and discarding the rest. The greedy search risks premature convergence and resource wastage. Instead, REvolve uses genetic programming to evolve a population of individuals, thus mitigating such risks at no extra computational cost.
Evolutionary Algorithms with LLMs. Implementing evolutionary algorithms with LLMs has been explored in recent works in the context of neural architecture search [34], program search [41], prompt engineering [11, 14], as well as morphology design [27]. In this work, we propose a novel evolutionary framework for reward design in AD with LLMs as intelligent genetic operators.
5 Experiments
For our experiments, we used the AirSim simulator [48], a high-fidelity simulator for AD research. We used the GPT-4 Turbo (1106-preview) model as the reward designer $G$, and provided it with (a subset of) observation and action variables originating from the AirSim simulator [^2] To prevent potential plagiarism through context memory [5], we use unique environment variable names and deliberately avoid mentioning the AirSim simulator in inputs to GPT-4., as illustrated in Figure 1.
5.1 Baselines
Human Driving. To benchmark the learned policies against human driving, we recruited participants to drive in a highly realistic setup, conjoining the AirSim (simulated) environment with a Logitech driving simulator. We collected data from 3 participants, each with 10 test runs. Participants were allowed trial runs to familiarize themselves with the setup.
Expert Designed. We used an expert reward function based on best practices established in the AD literature [44, 52]. The total reward is calculated through a weighted aggregation of all the components, with the weights optimized via an exhaustive grid search conducted within the AirSim simulator in order to find the reward function that produces the best policy. Details on the expert-designed reward function can be found in Appendix C.
Eureka. Closest to our method is Eureka [30], which creates a population of reward functions per iteration and greedily iterates on the best individual based on feedback. We employed the human fitness function to refine the best individual with natural language feedback, akin to REvolve.
Additionally, to emphasize the importance of human fitness function, in general, and human feedback, in particular, we explored an automatic feedback mechanism originally proposed in Eureka. Here, the (string) feedback is composed of the statistics of all reward components tracked from intermediate training checkpoints. e.g., "collision_penalty: [-1, -0.8, -0.4, …]". Ideally, a component is deemed beneficial if its statistics consistently grow/fall during training and potentially less useful if they are constant. Moreover, the fitness score is computed using a manually designed fitness function, which evaluates the policy on key performance indicators like collisions (Boolean), speed (real), and distance covered (real). We refer to Appendix B.2 for further details. We call these variants of Eureka and REvolve, Eureka Auto and REvolve Auto, respectively. Lastly, we also included T2R [61] in our comparison, which outputs a single reward function and iteratively refines it using feedback (T2R Auto).
Figure 4: Figure compares REvolve with baselines using manually designed fitness scores. It can be observed that REvolve consistently improves over generations, achieving its highest fitness score $\sigma=0.84$. The upper limit is marked by Human Driving at $\sigma=0.90$ while the Expert Designed fitness score is fixed at $\sigma=0.65$.
Figure 5: Figure compares policy training with REvolve and Expert Designed reward functions, on episodic steps. Policy training converges at a higher value for REvolve signifying a higher success rate. This is true for REvolve and REvolve Auto, with the former outperforming the latter.
5.2 Training Details
The observation of a policy includes both visual (egocentric camera view) and sensor data (e.g. yaw, pitch, linear velocity). The action space $A$ consists of steering and throttling controls, with a size $|A|=66$. The objective is to navigate an urban environment for 1000 consecutive time steps. A successful termination of the episode happens when the agent meets the objective. Failures like collisions or breaking speed thresholds will also terminate the episode. The specific details of the RL setup is outlined in Appendix A.
REvolve Hyperparameters: We set the number of generations to $N=5$ and individuals per generation to $K=16$, similar to Eureka. The number of sub-populations was set to $I=13$, and the mutation probability was set to $p_{m}=0.5$. For each generation, $16$ policies were trained (corresponding to each reward function) over $5\times 10^{5}$ training steps, on 16 NVIDIA A100 GPU (40GB). The training time for each policy was $\approx 50$ hours. Due to the extensive computational demands and significant training times, we report on a single random seed for both REvolve and Eureka baselines. Regarding human feedback, we engaged 10 human evaluators per generation, each assessing 20 pairs of videos. No personal data was collected.
5.3 Experimental Questions & Results
How does REvolve fair against the baselines on manually designed fitness scores?
As shown in Figure 4, REvolve exhibits continuous improvement across successive generations, ultimately achieving a higher fitness score on the manually designed fitness scale. Furthermore, both REvolve ($\sigma=0.84$) and REvolve Auto ($\sigma=0.81$) outperform their Eureka counterparts ($\sigma=0.72$ and $\sigma=0.70$, respectively), indicating that evolutionary search strategies yield better results than greedy search methods [^3] Note that these results are not displayed on the human fitness scale because the Elo ranking system evaluates individuals relative to one another. Therefore, comparisons between individuals from different baselines are only feasible if we directly pair them against one another.. We can also see that all methods but T2R Auto outperform the Expert Designed reward function ($\sigma=0.65$). Furthermore, we also observe the benefit of providing human feedback (REvolve: $\sigma=0.84$, Eureka: $\sigma=0.72$), as both REvolve and Eureka with provided human feedback outperforms their automatic feedback counterparts (REvolve Auto: $\sigma=0.81$, Eureka Auto: $\sigma=0.70$). None of the methods, however, surpassed the Human Driving baseline with $\sigma=0.90$, calculated by averaging the fitness scores over 30 runs (from 3 participants).
Do human fitness scores induce human-aligned behavior?
We further compared REvolve and Eureka on human evaluations. To this end, we generated rollouts from the best performing policies ($\pi^{\ast}$) from REvolve and Eureka, along with Human Driving. We paired them against each other and used the human feedback interface (as seen in Figure 1) for collecting human preference data. Collected human preference data was subsequently passed through our Elo rating system to rank each behavior. As shown in Table 1, REvolve policies secure second and third ranks, surpassed only by Human Driving.
We show the best (i.e. fittest) REvolve-designed reward functions $R^{\ast}$ in Appendix F.
Do REvolve reward functions lead to more episodic steps?
Episodic steps denote the number of actions taken by an agent within an episode before termination– either by reaching the episode length ($10^{3}$ steps) or by taking a non-recoverable action (e.g. crashing). In Figure 5 we see that REvolve-designed rewards converge to a higher number of episodic steps compared to expert-designed rewards, signifying a higher rate of successful actions per episode.
Do REvolve-designed reward functions generalize to new environments?
Aside from our standard environment (seen in Figure 6), we assessed the generalizability of the best REvolve-designed reward functions, $R^{\ast}$ in two new AirSim environments. Specifically, we trained policies from scratch using $R^{\ast}$ in two new scenarios: (Env 1) featuring lanes and a completely altered landscape while maintaining similar traffic conditions as our standard environment, and (Env 2) characterized by increased traffic with multiple cars actively maneuvering. From Table 1, we observe that in both environments, REvolve outperforms expert-designed rewards. However, the generalizability was lower for Env 2. This issue may stem from the temporal processing limitations of the RL policy’s visual module, which sometimes fails to detect oncoming cars. Implementing more advanced models could enhance detection capabilities [18]. Figure 7 in the Appendix A provides visual representations of these environments.
Table 1: The left table shows the best policies $\pi^{\ast}$ from each framework ranked based on human preferences, alongside Human Driving. We use the Elo rating system to map preferences to scores. All policies are initially assigned an Elo score of 1500. The right table shows the generalization scores in different environments. Here Fitness denotes manually designed fitness scores (mean ± std). Env 1 and Env 2 have 1500 and 1000 episodic steps, respectively.
Rank Baseline Elo Score
1 Human Driving 1595
2 REvolve 1534
3 REvolve Auto 1518
4 Eureka 1453
5 Eureka Auto 1401
Env Baseline Fitness Episodic Steps
1 Expert Designed 0.71 ± 0.03 710 ± 199
Human Driving 0.95 ± 0.02 1500 ± 0
REvolve 0.86 ± 0.03 1146 ± 297
2 Expert Designed 0.48 ± 0.02 71 ± 22
Human Driving 0.75 ± 0.04 1000 ± 0
REvolve 0.68 ± 0.04 135 ± 31
5.4 Limitations.
As stated in § 5.2, due to its (deep) RL nature, training a policy given a reward function is rather computationally expensive and time-consuming. While the computational cost of REvolve is the same as for Eureka – with no resources wasted on discarding individuals – REvolve does incur a slightly higher storage cost due to the maintenance of the reward database. Another limitation of REvolve is its reliance on the closed source GPT-4 model. This was a necessary choice as existing works have shown that a considerable gap exists in RD abilities between GPT-4 and open-source LLMs [30, 61]. With rapid advancements made by the open-source community, we believe this gap will be bridged in the near future.
6 Conclusion & Future Work
We formulated RDP as a search problem using evolutionary algorithms for searching through the space of reward functions (§ 2). With REvolve, we proposed a novel evolutionary framework that integrates LLMs as intelligent operators refining solutions through natural language feedback (§ 3). While we demonstrate our framework on the challenging problem of AD (§ 5), we also note that the generality of the formulation allows for broad applicability, especially in better alignment of deep learning models with human values and objectives. Unfortunately, the proposed framework can also be exploited by bad actors.
A key future question for REvolve is its real-world applicability. One promising approach involves integrating REvolve with Sim2Real algorithms [1, 20] by first learning a policy in simulation, using REvolve-designed rewards, and then transferring it to the real world. Recently, Ma et al. [31] have demonstrated that such an approach is viable for Eureka by tackling a quadruped balancing task on a physical robot. Moreover, our framework opens up opportunities for future research in Human-AI interaction. In particular, exploring the impact of more open-ended feedback from various perspectives, such as egocentric or third-person views.
Acknowledgement
This work was fully supported by the Knut and Alice Wallenberg foundation via the Wallenberg AI, Autonomous Systems and Software Program (WASP) and the Wallenberg scholar grant. We also acknowledge the Berzelius cluster at the National Supercomputer Centre, Sweden, for granting access to high-performance computing resources under projects "Berzelius-2024-178" and "Berzelius-2023-348". Finally, we thank Luc De Raedt for his valuable feedback in refining the manuscript.
Appendix for "REvolve: Reward Evolution with
Large Language Models for Autonomous Driving"
The appendix is organized as follows: 1. RL Preliminaries § A where we elaborate on the reinforcement learning framework. 2. Fitness Function § B where we outline the fitness score computation from human feedback. Additionally, we outline the automatic feedback approach from Eureka. 3. Expert Designed Reward Function § C which provide details of the expert designed reward and grid search parameters. 4. Prompts § D: System and operator (mutation, crossover) prompts for GPT-4 for REvolve and REvolve Auto 5. Human Feedback Examples § E to show how human feedback allows GPT-4 to refine the reward function individual. 6. Additional Results § F provides best reward functions generated from REvolve, and some additional results.
We also include our code and demonstrations of policies trained on best REvolve-designed reward functions, in the supplementary material.
Appendix A RL Preliminaries
A.1 Setup
Simulation Environment. The experiments are conducted in the neighborhood environment, a high-fidelity simulator developed for autonomous vehicle research [48]. This urban simulation includes static features such as parked cars and dynamic obstacles like cars and moving animals, enabling a thorough evaluation of the RL agent in varied scenarios.
Figure 6: AirSim simulator environment used for our AD experiments. The egocentric camera view displayed on the bottom left serves as the observations for the AD agent.
Figure 7: Airsim generalization environments used for assessing the generalizability of REvolve-designed reward functions. The left and right pictures visualize Env 1 and Env 2 respectively.
Observations. The observation space for the policy includes a combination of visual and sensor data. The visual input consists of images concatenated along the channel axis, resulting in dimensions of $100\times 256\times 12$. The sensor data includes readings from yaw, pitch, linear velocity in the x and y directions, speed, and steering angle. This structure allows the network to perceive temporal dynamics by integrating sequential frames into a single input tensor.
The input shape is $(100,256,12)$, and it outputs $66$ discrete Q-values. To capture temporal dynamics, frame stacking aggregates four consecutive frames $[fr_{1},fr_{2},fr_{3},fr_{4}]$ along the channel dimension into a single state. For subsequent states, the oldest frame is replaced by a new one, resulting in an updated stack $[fr_{2},fr_{3},fr_{4},fr_{\text{new}}]$.
Action Space. The action space for the autonomous car is a 2D vector – the first dimension represents the discrete steering angles in the range $[-32^{\circ},32^{\circ}]$ in $2^{\circ}$ increments, the second dimension is a throttling option $\{0,1\}$. This results in 66 possible actions.
Objective. The agent’s task is to navigate an urban environment for 1000 consecutive time steps, maintaining safe driving behavior by balancing speed and avoiding collisions.
Termination. The episode successfully terminates when the agent achieves the objective. Alternatively, collisions or exceeding speed limits will also prematurely end the episode.
A.2 Algorithm
We adopt the Clipped Double Deep Q-Learning approach [12], a refined version of Double Q-Learning [56]. This method uses two separate neural networks to reduce overestimation bias by calculating the minimum of the two estimated Q-values, thus providing more reliable and stable training outcomes.
In addition to Clipped Double Q-Learning, our framework integrates the Dueling Architecture [60], which improves the estimation of state values by decomposing the Q-function into two streams: one that estimates the state value and another that calculates the advantages of each action. Such structure allows the model to learn which states are valuable without the need to learn each action effect, for each state.
Furthermore, we utilize Prioritized Experience Replay (PER) [47] to focus on significant learning experiences. PER prioritizes transitions with higher Temporal Difference errors for replay, which accelerates learning and increases the efficiency of the memory used during training. Table 2 lists the RL model hyperparameters. Finally, inspired by foundational research, we have adapted our architecture to incorporate vision-based methods and the Dueling Architecture from [33] and [60], enhancing temporal dynamics and action-value precision.
Table 2: DDQN Hyperparameters
Parameter Value
Learning Rate 0.00025
Optimizer Adam
Gamma 0.99
Epsilon Initial 1
Epsilon Min 0.01
Epsilon Decay $1\times 10^{-3}$
Batch Size 32
Tau 0.0075
Replay Buffer Size 50000
Alpha (PER) 0.9
Frequency Steps Update 5
Beta 0.4
Action Space Steering (Simulator) [-0.8, 0.8]
Action Space Throttle (Simulator) [0,1]
Action Space (Real steering) [-32°, 32°]
Image Res. $100\times 256\times 12$
Conv Layers 5 (16-256 Filters)
Dense Units after Flatten $256->[128\times 2]->66$
Sensor Input Dimensions 6 (Yaw, Pitch, $V_{x}$, $V_{y}$, Speed, Steering)
Appendix B Fitness Function
In this section, we provide details of (1) REvolve fitness function which is based on human evaluations; (2) the automatic feedback approach proposed in Eureka.
B.1 REvolve Fitness
As state in § 3.3, REvolve employs a user interface to gather human feedback on behavior $t$ sampled from a trained policy. Human evaluators review pairs of videos, submitting their preferences. A fitness function then maps this data to a real-valued fitness score and a natural language feedback i.e. $(\sigma,\lambda)\coloneqq F(t)$. We start by introducing the Elo rating system that maps the preference data to a fitness score. For two reward function individuals $A$ and $B$ with fitness scores $\sigma_{A}$ and $\sigma_{B}$ respectively, the expected scores $E_{A}$ and $E_{B}$ are calculated using the logistic function:
$E_{A}=\frac{1}{1+10^{(\sigma_{B}-\sigma_{A})/400}}\;\;\;\;\;\;\;\;E_{B}=\frac{1}{1+10^{(\sigma_{A}-\sigma_{B})/400}}$
We then update the individual fitness based on the expected scores.
$\sigma_{A}=\sigma_{A}+K\cdot(f(A,B)-E_{A})$
[…]
$f(\text{X, Y})=\begin{cases}1&\text{if the X wins against Y},\\ 0&\text{if the X loses to Y},\\ 0.5&\text{if the match results in a tie}.\end{cases}$
Here, $K=32$ determines the magnitude of fitness change. Intuitively, if higher-ranked individuals lose to those with lower rankings, they incur a more significant penalty, while the reverse scenario leads to a greater increase for the lower-ranked individual. Conversely, when two similarly ranked individuals compete, the change in fitness scores resulting from a win or loss is smaller. All individuals start with an initial rating of 1500.
For the natural language feedback $\lambda$, the checkboxes selected by evaluators – indicating aspects of the behavior that are satisfactory or require improvement – are stitched together using a natural language (NL) template. This final string serves as input for GPT-4.
The process greatly simplifies the feedback interface, as participants only need to choose their preferred video from a pair, rather than providing an absolute rating. Additionally, this method reduces individual biases that could emerge from various factors, such as the need to adjust ratings at different stages of evolution – from the initial phase, where many policies underperform, to later stages, when performances improve. To overcome this, we compare individuals not only within the same generation but also across different generations, updating the fitness scores of all individuals at the end of each generation [^4] This approach is akin to comparing chess grandmasters Gary Kasparov and Magnus Carlsen, each a dominant player in his respective era. However, a direct comparison would require them to be matched in their primes; otherwise, the basis of comparison would be skewed and unjustified.. In direct rating-based systems, participants must repeatedly rate all generations, adjusting their scales each time. In contrast, our preference-based rating system allows us to efficiently utilize all existing preference data across generations to assign new fitness scores, thus requiring minimal data collection for the current generation.
B.2 Eureka Automatic Feedback
Manually Designed Fitness Function.
The fitness function is primarily designed to evaluate the performance of an autonomous vehicle by assessing its behavior in speed regulation, path following, and collision occurrences. It assesses how well the vehicle adheres to speed regulations and how accurately it follows its designated waypoints while penalizing collisions and any deviations from critical operational thresholds. The function is structured as follows:
$\text{TotalFitness}(collision,v,d)=\begin{cases}\text{collision penalty},&\text{if collision occurs}\\ \text{speed penalty},&\text{if }v<v_{\text{min\_limit}}\text{ or }v>v_{\text{max\_limit}}\\ \text{distance penalty},&\text{if }d>d_{\text{fail}}\\ \text{score}(v,d),&\text{otherwise}\end{cases}$ (2)
where $v_{\text{adj\_min}}=v_{\text{min}}-v_{\text{threshold}},\quad v_{\text{adj\_max}}=v_{\text{max}}+v_{\text{th}}$, with $v_{\text{th}}$ small threshold speed allowing for small deviations from the target speed range.
Here, penalties are assigned for: collision penalty$=-1$ for collisions; speed penalty is calculated based on the deviation from the adjusted range $[v_{\text{adjusted\_min}},v_{\text{adjusted\_max}}]$, decreasing as the speed moves further away from this range, with the speed score set to 0 if the speed is less than $v_{\text{min\_limit}}$ or greater than $v_{\text{max\_limit}}$; and distance penalty decreases as the vehicle moves further from the intended waypoint, with the distance score set to 0 if the deviation is above $d_{\text{fail}}$. Otherwise, the agent receives a $\text{score}(v,d)=\frac{\text{speed score}(v)+\text{distance score}(d)}{2}$.
The speed score is calculated as:
speed $\text{speed score}(v)=\begin{cases}1,&\text{if }v_{\text{adj\_min}}\leq v\leq v_{\text{max}}\\ \max(0,1-\frac{\min(|v-v_{\text{adj\_min}}|,|v-v_{\text{adj\_max}}|)}{v_{\text{adj\_max}}-v_{\text{adj\_min}}}),&\text{otherwise}\end{cases}$
The speed range is $[8.0,11.5]$ m/s. The distance score is calculated as:
distance score($\text{distance score}(d)=\begin{cases}1,&\text{if }d\leq d_{\text{max}}\\ \max(0,1-\frac{d-d_{\text{max}}}{d_{\text{max}}}),&\text{otherwise}\end{cases}$
with $d_{max}=0.5$m being the maximum tolerance for deviation from the waypoint. Both functions are shown in 8.
Figure 8: Visual representation of speed and distance scores.
In table 3, we summarize the fitness function components used for evaluating RL models.
Table 3: Fitness Function Hyperparameters Details
Hyperparameter Description Value
collision_penalty Collision penalty -1
$v_{\text{min}}$ Minimum speed target 9.0 m/s
$v_{\text{max}}$ Maximum speed target 10.5 m/s
$v_{\text{min\_limit}}$ Minimum penalty limit 2.5 m/s
$v_{\text{max\_limit}}$ Maximum penalty limit 15 m/s
$v_{\text{adjusted\_min}}$ Adjusted minimum speed 8 m/s
$v_{\text{adjusted\_max}}$ Adjusted maximum speed 11.5 m/s
$v_{{}_{th}}$ Threshold Speed 1 m/s
$d_{\text{max}}$ Waypoint deviation tolerance 0.5 m
$d_{\text{fail}}$ Maximum tolerance for deviation from the waypoint 4 m
Feedback.
Following Eureka, we track the statistics of all reward components at intermediate policy checkpoints throughout training. We refer the interested reader to the original paper [30].
Appendix C Expert Designed Rewards
Following the methodologies advocated in [44, 52], the total reward at each time step, $R_{\text{total}}$, is constructed from a composite of critical driving behaviors. This section describes three configurations of speed rewards, each designed to optimize the agent’s driving performance by emphasizing speed consistency, smooth navigation, and collision avoidance.
Human Engineered:
Utilizes a combination of position, speed, sensor, and smoothness rewards.
Total Reward:
$r_{\text{total}}^{HE}=c_{\text{pos}}\times r_{\text{pos}}+c_{\text{sm}}\times r_{\text{sm}}+c_{\text{speed}}\times r_{\text{speed}}+c_{\text{sensor}}\times r_{\text{sensor}}$
Position Reward:
Encourages the vehicle to stay close to the target position, rewarding proximity to the desired path or waypoint.
$r_{\text{pos}}=e^{-\delta\times((\text{min\_pos})^{2}-\beta)}$
Speed Reward:
Promotes maintaining an optimal speed. It penalizes the vehicle if it moves too slowly or too quickly, encouraging it to stay within a specified speed range.
$r_{\text{speed}}(\text{speed})=\begin{cases}-\frac{|\text{speed}-v_{\text{min}}|}{v_{\text{min}}}&\text{if speed}<v_{\text{min}}\\ -\frac{|\text{speed}-v_{\text{max}}|}{\text{speed}}&\text{if speed}>v_{\text{max}}\\ 1&\text{otherwise}\end{cases}$
Sensor Reward:
Encourages the vehicle to maintain a safe distance from obstacles using sensor data. If the distance measured by the sensor is less than a safe threshold, it applies a penalty. The sensor measures the distance from the vehicle to nearby front objects.
$r_{\text{sensor}}=\begin{cases}-\frac{(dis_{safe}-\text{distance})}{dis_{safe}}&\text{if distance}<dis_{safe}\\ 0.5&\text{otherwise}\end{cases}$
Smoothness Reward:
Penalizes abrupt changes in steering to encourage smooth driving. It uses the standard deviation of the last four steering actions to measure smoothness.
[…]
Table 4: Values and Corresponding Fitness Scores
c_pos c_sm c_speed c_sensor Fitness Score
0.85 0.46 0.5 0.53 0.496
0.25 0.25 0.25 0.25 0.603
0.81 0.83 0.33 0.13 0.606
0.35 0.67 0.7 0.67 0.576
0.57 0.26 0.57 0.27 0.396
0.22 0.16 0.24 0.85 0.556
0.67 0.3 0.78 0.77 0.549
0.45 0.11 0.27 0.24 0.256
0.72 0.93 0.74 0.65 0.415
0.63 0.4 0.39 0.78 0.284
0.44 0.76 0.98 0.48 0.653
0.28 0.64 0.2 0.92 0.602
0.12 0.52 0.91 0.41 0.592
0.94 0.86 0.85 0.2 0.571
0.51 1.0 0.64 0.34 0.630
0.98 0.56 0.12 0.95 0.600
0.88 0.38 0.46 0.59 0.591
Table 5: Reward Components Hyperparameters Details
Hyperparameter Description Value
$\delta$ Position reward decay factor 0.2
$\beta$ Position reward bias 0.25
$v_{\text{min}}$ Minimum target speed 9 m/s
$v_{\text{max}}$ Maximum target speed 10.5 m/s
$dis_{safe}$ Safe distance threshold 6 m
$\gamma$ Smoothness penalty coefficient 0.5
Human Engineered Tuned:
We used grid search to optimize the values of the hyperparameters $c_{\text{pos}},c_{\text{sm}},c_{\text{speed}},c_{\text{sensor}}$.
Total Reward:
$r_{\text{total}}^{HET}=c_{\text{pos}}^{opt}\times r_{\text{pos}}+c_{\text{sm}}^{opt}\times r_{\text{sm}}+c_{\text{speed}}^{opt}\times r_{\text{speed}}+c_{\text{sensor}}^{opt}\times r_{\text{sensor}}$
These configurations allow for tailored training of autonomous driving agents, emphasizing different aspects of driving performance based on specific needs and expected driving conditions. All configurations use weights of $c_{\text{pos}},c_{\text{sm}},c_{\text{speed}},c_{\text{sensor}}=0.25$, except for the optimized version which uses customized weights.
Appendix D Prompts
Here, we present the different prompts used for GPT-4. It includes the following:
System Prompt In Box D that outlines the task description and expected input from the environment, complete with formatting tips to help guide GPT-4 in generating suitable reward functions.
Abstracted Environment Variables In Box D that provides a list of (abstracted) observation and action variables in the environment, formatted as a Python class. This informs GPT-4 of variables it can use in reward design.
Operator Prompts like crossover (Box D) and mutation (Box D). Additionally, we provide prompts for REvolve Auto baseline (Boxes D D).
Box 1: System Prompt You are a reward engineer trying to write a reward function to create an autonomous car agent to learn to navigate the environment. Your goal is to write a reward function for the environment that will help the agent learn the task described in text.
Task description:
The objective is to create an autonomous car agent to learn to navigate the environment. Specifically, the performance metric for the learning agent is to successfully drive without crashing, while maintaining smooth driving behavior and a speed between 9.0-10.5 m/s. The episode terminates successfully if the agent drives for 1000 consecutive time steps without crashing. Otherwise, the episode terminates in failure. The controlled actions are the steering and the speed. The car must avoid staying inactive and stay as close to the center of the road as possible to ensure optimal performance.
Your reward function should use useful variables from the environment as inputs. An example reward function signature can be: ```python def compute_reward(object_pos: type, goal_pos: type) -> Tuple[type, Dict[str, type]]: ... return reward, {} ``` The output of the reward function should consist of two items: 1. The total reward, 2. A dictionary of each individual reward component. The code output should be formatted as a Python code string: "“‘python … “‘". 1. You may find it helpful to normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward and/or its reward components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable and assume variables that are not available. 5. Your output must always be the def function code with the output being the total reward. 6. Make sure you don’t give contradictory reward components.
Box 2: Abstracted Environment Variables (as a Python class) from scipy.spatial.transform import Rotation as R import numpy as np class Env: def __init__(self): self.client = CarClient() action = [CarControls().steering, CarControls().throttle] def state(self): position_info = self.client.getCarState().kinematics_estimated curr_x = position_info.position.x_val curr_y = position_info.position.y_val yaw, pitch = self.get_orientation(position_info) linear_velocity = position_info.linear_velocity vx, vy = linear_velocity.x_val, linear_velocity.y_val angular_velocity_x, angular_velocity_y, angular_velocity_z = position_info.kinematics_estimated.angular_velocity.x_val, position_info.kinematics_estimated.angular_velocity.y_val, position_info.kinematics_estimated.angular_velocity.z_val speed = position_info.speed collision = self.client.simGetCollisionInfo().has_collided action_list = [steering_value1, steering_value2, steering_value3, steering_value4] min_pos = get_min_pos(self, position_info) total_step_counter = 0 episode_step_counter = 0 distance = 20 # default max range fo sensor is 20 def get_orientation(self, position_info): quaternion = np.array([position_info.orientation.w_val, position_info.orientation.x_val, position_info.orientation.y_val, position_info.orientation.z_val]) euler_angles = R.from_quat(quaternion).as_euler(’zyx’, degrees=True) return euler_angles[0], euler_angles[1] def get_min_pos(self, position_info): min_dis = float(’inf’) for i in data: stx, sty = i[0], i[1] dis = eucl_dis(stx, sty, curr_x, curr_y) if dis < min_dis: min_dis = dis return min_dis def calculate_front_distance(self): distance_data = self.client.getDistanceSensorData(vehicle_name="Car") distance = distance_data.distance # max sensor distance is 20
Box 3: REvolve Crossover Prompt You are given a set of reward functions used to train autonomous driving agents. The reward functions have fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. For each reward function, we collected human feedback on what aspects of the agent’s driving are satisfactory and what aspects need improvement. Additionally, we tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed.  <EXAMPLES> Your task: Combine high-performing reward components from different reward functions based on human feedback and the data tracked, to enhance the agent’s driving capability further. Clearly explain how you intend to combine them and your rationale behind improving the driving performance. Write the combined reward function code. Output of the reward function should consist of: 1: The total reward. 2: A dictionary of each individual reward component. Code output should be formatted as a Python code string: ```python``` Tips for reward function design: 1. Normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward or its components. 2. If transforming a reward component, introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. Each transformed reward component should have its own temperature variable and you must carefully set its value based on its effect in the overall reward. 3. Ensure the type of each input variable is correctly specified. 4. The reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable. 5. Always output the def function code with the output being the total reward. 6. Avoid contradictory reward components. 7. Ensure that the penalties for collision and inactivity are given precedence over other reward components. This means that in the event of a collision or the car staying still, the total reward for that time step should primarily reflect the penalty for these specific conditions, potentially negating other positive rewards accrued from different aspects of the driving performance.
Box 4: REvolve Mutation Prompt You are given a reward function used to train autonomous driving agents. The reward functions has a fitness scores that reflects the agent’s driving performance – higher scores indicate superior driving behavior. We also collected human feedback on what aspects of the agent’s driving are satisfactory and what aspects need improvement. Additionally, we tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed.
<EXAMPLES>
Your task is to iterate on the reward function by mutating a single component to enhance the agent’s driving capability further. Some helpful tips for analyzing the reward components: If the values for a certain reward component are nearly identical throughout, then this means the training is not able to optimize this component as it is written. You may consider:
1. Rescaling it to a proper range or the value of its temperature parameter 2. Re-writing the reward component 3. Discarding the reward component The mutation process involves the following steps: 1. First, select a reward component based on the given guidelines, human feedback, and the values of individual reward components. 2. Next, clearly explain how you intend to mutate the selected component and your rationale behind improving the driving performance. This could involve adjusting its scale, rewriting its formula, or other modifications. 3. Finally, write the mutated reward function code.
The output of the reward function should consist of two items: 1. The total reward, 2. A dictionary of each individual reward component. The code output should be formatted as a Python code string: ```python``` Some helpful tips for writing the reward function code: 1. You may find it helpful to normalize the reward to a fixed reward range like -1,1 or 0,1 by applying transformations like tf.exp() to the overall reward or its components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. Each transformed reward component should have its own temperature variable and you must carefully set its value based on its effect in the overall reward. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable.
Box 5: REvolve Auto Crossover Prompt You are given a set of reward functions used to train autonomous driving agents. The reward functions have fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. We also tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed.
<EXAMPLES>
Your task: Combine high-performing reward components from different reward functions to enhance the agent’s driving capability further. Clearly explain how you intend to combine them and your rationale behind improving the driving performance. Write the combined reward function code.
Output of the reward function should consist of: 1: The total reward. 2: A dictionary of each individual reward component.
Code output should be formatted as a Python code string: ```python```
Tips for reward function design: 1. Normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward and/or its reward components. 2. If transforming a reward component, introduce a temperature parameter inside the transformation function. 3. Ensure the type of each input variable is correctly specified. 4. The reward’s code input variables must contain only attributes of the environment. 5. Always output the def function code with the output being the total reward. 6. Avoid contradictory reward components.
Box 6: REvolve Auto Mutation Prompt You are given a reward function used to train autonomous driving agents. The reward function has fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. We also tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, and minimum values observed.
<EXAMPLES>
Your task is to iterate on the reward function by mutating a single component to enhance the agent’s driving capability further. Some helpful tips for analyzing the reward components: If the values for a certain reward component are nearly identical throughout, then this means the training is not able to optimize this component as it is written. You may consider:
1. Rescaling it to a proper range or the value of its temperature parameter 2. Re-writing the reward component 3. Discarding the reward component
The mutation process involves the following steps:
1. First, select a reward component based on the given guidelines. 2. Next, clearly explain how you intend to mutate the selected component and your rationale behind improving the driving performance. This could involve adjusting its scale, rewriting its formula, or other modifications. 3. Finally, write the mutated reward function code.
The output of the reward function should consist of two items:
1. The total reward, 2. A dictionary of each individual reward component.
The code output should be formatted as a Python code string: ```python``` Some helpful tips for writing the reward function code:
1. You may find it helpful to normalize the reward to a fixed reward range like -1,1 or 0,1 by applying transformations like tf.exp() to the overall reward or its components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and not an input variable. Each transformed reward component should have its own temperature variable, and you must carefully set its value based on its effect on the overall reward. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable.
Appendix E Human Feedback Examples
Box 7: REvolve Mutation Prompt Human feedback on the autonomous driving agent’s performance: Positive: collision avoidance, lane keeping, managing consistent speed, intersection and turn handling.
Negative: Smooth Steering.
Based on the human feedback and analysis of the reward components, we can identify that the "smooth driving bonus" needs modification. The feedback indicates issues with smooth steering, which suggests that the existing smoothness component isn’t effectively promoting the desired steering behavior. Given that the human feedback specifically pointed out deficiencies in this area, it suggests the measure (variance of steering actions) might not be adequately sensitive or is scaled improperly.
Intended Mutation: The mutation will focus on revising the "smooth driving bonus" to better capture and reward smooth steering. The current method uses the standard deviation of steering actions to compute a penalty, which may not effectively discriminate between slightly varying and highly erratic steering actions. A more responsive approach might involve squaring the deviation before scaling, which penalizes larger variations more severely and provides a sharper gradient for the learning algorithm to optimize against.
Rationale By squaring the deviation, small steering adjustments are less penalized, allowing for natural minor corrections while driving. Larger, more erratic movements result in significantly higher penalties, pushing the training to favor smoother transitions and more stable steering behaviors.
def old_compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: ... max_steering_variance = 0.1 smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor / max_steering_variance)) ... def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: ... # Updated mutation for smooth driving bonus max_steering_variance = 0.1 smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor**2 / max_steering_variance**2)) ...
Appendix F Additional Results
In Box F we showcase the best performing reward (fitness score of $0.84$) In Box F we showcase the best performing reward (fitness score of $0.81$)
Box 8: Best reward function from REvolve def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, min_pos: float, action_list: List[float]) -> Tuple[float, Dict[str, float]]: # Parameters to tweak the importance of different reward components collision_penalty = -100 inactivity_penalty = -10 speed_reward_weight = 2.0 position_reward_weight = 1.0 smoothness_reward_weight = 0.5 # Adjusted temperature parameters for score transformation speed_temp = 0.5 # Increased temperature for speed_reward position_temp = 0.1 smoothness_temp = 0.1 reward_components = { "collision_penalty": 0, "inactivity_penalty": 0, "speed_reward": 0, "position_reward": 0, "smoothness_reward": 0 } # Penalize for collision if collision: reward_components["collision_penalty"] = collision_penalty # Penalize for inactivity (speed too close to zero) if speed < 4.5: # Adjusted threshold for inactivity reward_components["inactivity_penalty"] = inactivity_penalty # Reward for maintaining an optimal speed range if 9.0 <= speed <= 10.5: speed_score = 1 - np.abs(speed - 9.75) / 1.75 # Centering and normalizing around the average of optimal speed range else: speed_score = -1 # Use a sigmoid function to smoothly transform the speed score and constrain it reward_components["speed_reward"] = speed_reward_weight * (1 / (1 + np.exp(-speed_score / speed_temp))) # Reward for being close to the center of the road position_score = np.exp(-min_pos / position_temp) reward_components["position_reward"] = position_reward_weight * position_score # Reward for smooth driving (small variations in consecutive steering actions) steering_smoothness = -np.std(action_list) reward_components["smoothness_reward"] = smoothness_reward_weight * np.exp(steering_smoothness / smoothness_temp) # Calculate total reward, giving precedence to penalties total_reward = 0 if reward_components["collision_penalty"] < 0: total_reward = reward_components["collision_penalty"] elif reward_components["inactivity_penalty"] < 0: total_reward = reward_components["inactivity_penalty"] else: total_reward = sum(reward_components.values()) # Ensure the total reward is within a reasonable range total_reward = np.clip(total_reward, -1, 1) return total_reward, reward_components
Box 9: Best reward function from REvolve Auto def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: # Define temperature parameters for reward transformations temp_collision = 5.0 temp_inactivity = 10.0 temp_centering = 2.0 temp_speed = 2.0 temp_smoothness = 1.0 temp_proximity = 2.0 # Initialize the total reward and reward components dictionary reward_components = { ’collision_penalty’: 0, ’inactivity_penalty’: 0, ’lane_centering_bonus’: 0, ’speed_regulation_bonus’: 0, ’smooth_driving_bonus’: 0, ’front_distance_bonus’: 0 } # Penalty for collisions reward_components[’collision_penalty’] = np.exp(-temp_collision) if collision else 0 # Penalty for inactivity inactivity_threshold = 4.5 if speed < inactivity_threshold and not collision: reward_components[’inactivity_penalty’] = np.exp(-temp_inactivity * (inactivity_threshold - speed)) # Lane centering bonus max_distance_from_center = 2.0 # adaptable based on road width reward_components[’lane_centering_bonus’] = np.exp(-temp_centering * min_pos) # Speed regulation bonus ideal_speed = (9.0 + 10.5) / 2 speed_range = 10.5 - 9.0 reward_components[’speed_regulation_bonus’] = np.exp(-temp_speed * (abs(speed - ideal_speed) / speed_range)) # Smooth driving bonus max_steering_variance = 0.1 # configured to the acceptable steering variance for full bonus smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor / max_steering_variance)) # Front distance bonus max_distance_view = 20 reward_components[’front_distance_bonus’] = np.clip(distance / max_distance_view, 0, 1) # Sum the individual rewards for the total reward total_reward = sum(reward_components.values()) # Ensure penalties take precedence if collision or speed < inactivity_threshold: total_reward = -1 # Apply max negative reward for collision or inactivity # Normalize the total reward to lie between -1 and 1 total_reward = np.clip(total_reward, -1, 1) return total_reward, reward_components
Figure 9: Compares REvolve with baselines using manually designed fitness scores (average score across populations). Again REvolve consistently improves over generations, achieving a higher average fitness score $\sigma=0.67$.
Figure 10: Compares policy training with REvolve and Expert designed reward functions, on Episodic Steps from the start of the training.
Figure 11: Shows the total Episodic Steps during training for Env 1 scenario.
Figure 12: Shows the total Episodic Steps during training for Env 2 scenario.